40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 10558     Entry time: Wed Oct 1 19:40:46 2014
Author: ericq 
Type: Update 
Category: LSC 
Subject: Arms IR aligned 
Summary:
  • Beamsplitter was put into MC refl path.
  • HWP was rotated to maximize power into PMC. 
  • MC autolocker locked, small alignment tweak led to WFS taking over
  • Light present on REFL, AS and POP!
  • After small adjustments to TTs and ETMY, locked Yarm with AS55, ran ASS. 
  • Adjusted AS camera and RFPD alignment for ASS'd AS beam. 
  • Left arm locked on AS55, aligned new POY beam onto POY11. Centered ITMY oplev while I was there. 
  • After adjusting digital POY11 demod angle with an excitation into ETMY, arms were POX/POY locked and ASS'ed.
  • PRM and SRM eyeball aligned

The IFO is ready for 3F DRMI comissioning 

ELOG V3.1.3-